# Channel Estimation for Beyond Diagonal RIS via Tensor Decomposition

### v1

@@ IV Tensor signal modeling
In this section, we recast the signal model for the received pilots using a tensor decomposition approach. In correspondence with the background material presented in Section II, we formulate the tensor signal models by starting from the general (fully-connected) case for presentation convenience. Then, we discuss the tensor signal formulation for the group-connected case, which is of practical interest due to its lower implementation complexity. The main expressions derived in this section will be exploited later in the derivation of the proposed channel estimation methods.
IV-A General fully-connected case
@@ IV-B Group-connected case
Let us now consider the group-connected architecture. Starting from the signal model in (14), and ommitting the noise term, we have
@@
Group-connected architectures are commonly adopted due to implementation complexity. In this case, the BD-RIS training tensor is “sparse” due to its block-diagonal structure. Figure 2 illustrates the decomposition of received pilot signals in tensor form, corresponding to a Tucker-2 decomposition composed of $Q$ blocks. This decomposition can also be viewed as a special block-term decomposition (BTD), more specifically, also referred to as a “type-2 BTD”, which represents a tensor into a sum of rank-$(\bar{N},\bar{N},\cdot)$ tensor blocks [24, 28]. From this figure, we can see that the assumption of a group-connected architecture implies a BD-RIS training tensor having a “block-diagonal” structure, the $k$-th frontal slice of which corresponds to the BD-RIS training matrix associated with the $k$-th block. Note that a fully-connected BD-RIS architecture is equivalent to having a fully dense BD-RIS training tensor in Figure 2, which does not have zero off-diagonal blocks. Although we focus on the group-connected case based on (28), any (sub-connected) architecture can be captured by the general tensor representation in (20), the difference being in the structure of the BD-RIS training tensor $\bm{\mathcal{S}}$.
Similarly to (24)-(26), we construct the unfoldings of the $q$-th group BD-RIS tensor as
@@
Hence, for the group-connected case, equivalent expressions can be obtained for the unfoldings of the received pilot tensor. The corresponding expressions for the 1-mode and 2-mode matrix unfoldings of the received pilot tensor can be obtained by rewriting (21) and (22) as a sum of $Q$ blocks
$\big{[}\bm{\mathcal{Y}}\big{]}_{\left(1\right)}=\bm{G}\cdot\underbrace{\text{blkdiag}\Big{(}[\bm{\mathcal{S}}^{(1)}]_{(1)},\ldots,[\bm{\mathcal{S}}^{(Q)}]_{(1)}\Big{)}}_{\bm{S}_{1}}\left[\begin{array}[]{c}\bm{I}_{K}\otimes\bm{H}^{\text{T}}_{1}\\ \vdots\\ \bm{I}_{K}\otimes\bm{H}^{\text{T}}_{Q}\end{array}\right]$ (35)
$\big{[}\bm{\mathcal{Y}}\big{]}_{\left(2\right)}=\bm{H}\cdot\underbrace{\text{blkdiag}\Big{(}[\bm{\mathcal{S}}^{(1)}]_{(2)},\ldots,[\bm{\mathcal{S}}^{(Q)}]_{(2)}\Big{)}}_{\bm{S}_{2}}\left[\begin{array}[]{c}\bm{I}_{K}\otimes\bm{G}^{\text{T}}_{1}\\ \vdots\\ \bm{I}_{K}\otimes\bm{G}^{\text{T}}_{Q}\end{array}\right]$ (39)
or, compactly,
$\big{[}\bm{\mathcal{Y}}\big{]}_{\left(1\right)}$ $=$ $\bm{G}\bm{S}_{1}\left(\bm{\overline{I}}\hskip 1.42271pt|\otimes|\hskip 1.42271pt\bm{H}\right)^{\text{T}},$ (40)
$\big{[}\bm{\mathcal{Y}}\big{]}_{\left(2\right)}$ $=$ $\bm{H}\bm{S}_{2}\left(\bm{\overline{I}}\hskip 1.42271pt|\otimes|\hskip 1.42271pt\bm{G}\right)^{\text{T}},$ (41)
where
$\bm{S}_{1}\doteq\text{blkdiag}\big{(}[\bm{\mathcal{S}}^{(1)}]_{(1)},\ldots,[\bm{\mathcal{S}}^{(Q)}]_{(1)}\big{)}\in\mathbb{C}^{\bar{N}Q\times\bar{N}KQ}$ (42)
$\bm{S}_{2}\doteq\text{blkdiag}\big{(}[\bm{\mathcal{S}}^{(1)}]_{(2)},\ldots,[\bm{\mathcal{S}}^{(Q)}]_{(2)}\big{)}\in\mathbb{C}^{\bar{N}Q\times\bar{N}KQ}$ (43)
with $\bm{\overline{I}}\doteq[\bm{I}_{K},\ldots,\bm{I}_{K}]\in\mathbb{C}^{K\times KQ}$. Likewise, the 3-mode matrix unfolding of the received pilot tensor can be obtained by rewriting (23) as a sum of $Q$ blocks
$\big{[}\bm{\mathcal{Y}}\big{]}_{\left(3\right)}$ $=$ $\underbrace{\Big{[}[\bm{\mathcal{S}}^{(1)}]_{(3)},\ldots,[\bm{\mathcal{S}}^{(Q)}]_{(3)}\Big{]}}_{\bm{S}_{3}}\left[\begin{array}[]{c}\bm{H}^{\text{T}}_{1}\otimes\bm{G}^{\text{T}}_{1}\\ \vdots\\ \bm{H}^{\text{T}}_{Q}\otimes\bm{G}^{\text{T}}_{Q}\end{array}\right]$ (47)
@@
Remark 1: The pairs of expressions (21)-(22) and (40)-(41) representing the 1-mode and 2-mode unfoldings are equivalent and interchangeable due to the following property
[…]
$\bm{S}_{n}=[\bm{\mathcal{S}}]_{(n)}\bm{P},\,n=1,2,$ (50)
where $\bm{P}\doteq[\bm{I}_{Q}\otimes\bm{e}^{(K)}_{1},\ldots,\bm{I}_{Q}\otimes\bm{e}^{(K)}_{K}]\otimes\bm{I}_{\bar{N}}$ is a permutation matrix of dimensions $\bar{N}QK\times\bar{N}KQ$ and $\bm{e}^{(K)}_{k}$ is a unit vector corresponding to the $k$-th column of $\bm{I}_{K}$. This equivalence implies that the 1-mode and 2-mode expressions in (40)-(41) are equivalent to those in (21)-(22) since the first ones are just permuted versions of the latter. Hence, both are valid for formulating the block Tucker alternating least squares (BTALS) channel estimation method. On the other hand, one should note that the 3-mode unfolding expression in (49) provides a more compact representation than the one in (23) when considering a group-connected architecture. This is due to the fact that working with the 3-mode unfolding of the BD-RIS tensor $\bm{\mathcal{S}}\in\mathbb{C}^{\bar{N}Q\times\bar{N}Q\times K}$ introduced unnecessary zeros due to its block structure (see Figure 2). $\bm{S}_{3}$ is instead a concatenation of the 3-mode unfoldings of $Q$ tensor blocks. Consequently, it is not a permuted version of $[\bm{\mathcal{S}}]_{(3)}$ (as opposed to the 1-mode and 2-mode unfoldings). Indeed, the first is a $K\times\bar{N}^{2}Q$ matrix, while the second is a longer and sparse $K\times\bar{N}^{2}Q^{2}$ matrix whose number of columns is increased by a factor of $Q$. The difference in the structures of $\{\bm{S}_{1},\bm{S}_{2}\}$ and $\bm{S}_{3}$ is a consequence of the “asymmetric” structure of the block Tucker model in (28) with respect to its third mode, which is an identity matrix and hence does not depend on the number $Q$ of blocks. Hence, in the next section, we will adopt the compact 3-mode expression in (49) to formulate the block Tucker Kronecker factorization (BTKF) algorithm for channel estimation.
@@
which corresponds to a PARAFAC tensor model for the received pilots. It is clear that the proposed BD-RIS Tucker model generalizes the diagonal RIS PARAFAC model. For further details on tensor modeling and algorithms for the diagonal RIS case, we refer the reader to [19].
@@ V Tensor-Based channel estimation methods
In this section, we formulate the proposed channel estimation methods by capitalizing on the tensor signal structures discussed in the previous section. Two algorithms are proposed to solve the channel estimation problem via decoupling the estimates of the involved channel matrices $\bm{G}$ and $\bm{H}$. The first one, BTKF, is a closed-form solution that extracts the channel estimates via a block Kronecker factorization problem, which boils down to solving a set of rank-one matrix approximation problems. The second one is termed BTALS, which directly estimates the channel matrices using an iterative procedure. We also discuss identifiability and uniqueness issues and their overall implications for system design. The design of the BD-RIS training tensor is a separate topic that will be discussed in the next section.
@@ V-A Block Tucker Kronecker factorization algorithm (BTKF) algorithm for decoupled channel estimation
where $\overline{\bm{Z}}^{(q)}\in\mathbb{C}^{M_{R}\bar{N}\times M_{T}\bar{N}}$ is a reshaped version of $\bm{Z}^{(q)}\in\mathbb{C}^{M_{R}M_{T}\times\bar{N}^{2}}$, $\bm{g}^{(q)}=\text{vec}\big{(}\bm{G}^{(q)}\big{)}\in\mathbb{C}^{M_{R}\bar{N}\times 1}$, and $\bm{h}^{(q)}=\text{vec}\big{(}\bm{H}^{(q)}\big{)}\in\mathbb{C}^{M_{T}\bar{N}\times 1}$. Defining the truncated singular value decomposition (SVD) of $\overline{\bm{Z}}^{{(q)}}=\bm{U}^{(q)}\bm{\Sigma}^{(q)}\bm{V}^{\text{H}(q)}$, the estimates of $\bm{h}^{(q)}$ and $\bm{g}^{(q)}$ corresponding respectively to the dominant left and right singular vectors $\bm{u}^{(q)}_{1}=\bm{U}^{(q)}_{.1}\in\mathbb{C}^{M_{R}\bar{N}\times 1}$ and $\bm{v}^{*(q)}_{1}=\bm{V}^{*(q)}_{.1}\in\mathbb{C}^{M_{T}\bar{N}\times 1}$
$\hat{\bm{g}}^{(q)}$ $=\bm{u}^{(q)}_{1},\quad\hat{\bm{G}}^{(q)}=\text{unvec}\big{(}\hat{\bm{g}}^{(q)}\big{)}_{M_{R}\times\bar{N}},$ (57)
$\hat{\bm{h}}^{(q)}$ $=\bm{v}^{*(q)}_{1},\quad\hat{\bm{H}}^{(q)}=\text{unvec}\big{(}\hat{\bm{h}}^{(q)}\big{)}_{M_{T}\times\bar{N}}.$ (58)
@@
A block diagram of the main steps of the BTFK algorithm is provided in Figure 3.
Figure 3: Block-diagram of the BTKF algorithm.
@@ V-B Block Tucker alternating least squares (BTALS) algorithm for decoupled channel estimation
It is clear that this is a nonlinear problem in the unknowns since the solution involves products of the coefficients of the channel matrices $\bm{H}$ and $\bm{G}$. However, it is known from tensor decomposition theory [30] that this problem can be efficiently solved through an alternating least squares (ALS)estimation algorithm [30, 31], a popular iterative method for fitting a tensor model thanks to its simplicity and monotonic convergence property. In our case, the algorithm yields decoupled estimates of the channel matrices by converting the bilinear problem in (61) into the following two simplest linear LS sub-problems. Specifically, to estimate the channels $\bm{H}$ and $\bm{G}$, we exploit the $1$-mode and $2$-mode unfoldings of the received pilots tensor $\bm{\mathcal{Y}}$, defined in (40)-(41), respectively. The bilinear alternating least squares (BALS) method solves the following two LS problems in an iterative way
[…]
@@
$\hat{\bm{G}}$ $=\big{[}\bm{\mathcal{Y}}\big{]}_{\left(1\right)}\left[\bm{S}_{1}\left(\bm{\overline{I}}\hskip 1.42271pt|\otimes|\hskip 1.42271pt\bm{H}\right)^{\text{T}}\right]^{+},$ (62)
$\hat{\bm{H}}$ $=\big{[}\bm{\mathcal{Y}}\big{]}_{\left(2\right)}\left[\bm{S}_{2}\left(\bm{\overline{I}}\hskip 1.42271pt|\otimes|\hskip 1.42271pt\bm{G}\right)^{\text{T}}\right]^{+}.$ (63)
Figure 4: Block-diagram of the BTALS algorithm.
The algorithm intertwines the LS estimates of $\bm{G}$ and $\bm{H}$. At each iteration, one matrix is updated based on a previously obtained estimate of the other matrix. This procedure is repeated until the convergence. Figure 4 illustrates the BTALS iterations process. An error measure computed at the end of the $i$-th iteration is given by $\epsilon_{(i)}\doteq\|\big{[}\bm{\mathcal{Y}}\big{]}_{\left(3\right)}-\hat{\bm{Y}}_{(i)}\|^{2}_{{\text{F}}}/\|\big{[}\bm{\mathcal{Y}}\big{]}_{\left(3\right)}\|^{2}_{\text{F}}$, where $\hat{\bm{Y}}_{(i)}\doteq[\bm{\mathcal{S}}]_{(3)}\big{(}\bm{\hat{H}}_{(i)}\otimes\bm{\hat{G}}_{(i)}\big{)}^{\text{T}}$. The convergence is declared when this error does not significantly change between two successive iterations, which implies that $|\epsilon_{(i)}-\epsilon_{(i-1)}|\leq\eta$, where $\eta$ is a convergence threshold. A summary of the steps of the proposed BALS algorithm is provided in Algorithm 2.
Such a bilinear ALS procedure may improve or maintain but cannot worsen the current fit, usually leading to a global minimum solution [15]. For simplicity, random initialization can be adopted. However, one may resort to enhanced initialization and acceleration procedures [15, 30]. Other initialization strategies that could be used to exploit the structure of the channel matrices are possible, although not advocated here. In our case, where the core tensor (represented by the BD-RIS structure) is known at the receiver, such enhancements are not required, and random initialization is enough to obtain satisfactory results.
V-C Discussion
From the above discussion, it is clear that the BTALS algorithm has less restrictive requirements on the training length $K$ than the BTKF algorithm. This comes from the fact that the latter only exploits the 3-mode unfolding of the received pilot tensor (as is the case with the LS method [5]), and channel decoupling is carried out after the LS filtering step, which is the bottleneck restricting the training length. On the other hand, the BTALS algorithm exploits the 1-mode and 2-mode unfoldings of the received pilot tensor to yield direct decoupled estimates of the two involved channels. The associated LS steps involve more relaxed requirements for the training length $K$ compared to BTKF.
@@ V-C Discussion
6:     Define the SVD of $\overline{\bm{Z}}^{(q)}=\bm{U}^{(q)}\bm{\Sigma}^{(q)}\bm{V}^{(q)\text{H}}$.
@@
$\hat{\bm{G}}^{(q)}=\text{unvec}\big{(}\bm{u}^{(q)}_{1}\big{)}_{M_{R}\times\bar{N}},\,\,\hat{\bm{H}}^{(q)}=\text{unvec}\big{(}\bm{v}^{*(q)}_{1}\big{)}_{M_{T}\times\bar{N}}.$
@@
$\bm{\hat{G}}_{(i)}=\big{[}\bm{\mathcal{Y}}\big{]}_{\left(1\right)}\left[\bm{S}_{1}\left(\bm{\overline{I}}\hskip 1.42271pt|\otimes|\hskip 1.42271pt\bm{H}_{(i-1)}\right)^{\text{T}}\right]^{+}.$
@@
$\bm{\hat{H}}_{(i)}=\big{[}\bm{\mathcal{Y}}\big{]}_{\left(2\right)}\left[\bm{S}_{2}\left(\bm{\overline{I}}\hskip 1.42271pt|\otimes|\hskip 1.42271pt\bm{G}_{(i)}\right)^{\text{T}}\right]^{+}.$
6:     Compute $\hat{\bm{Y}}_{(i)}=[\bm{\mathcal{S}}]_{(3)}\big{(}\bm{\hat{H}}_{(i)}\otimes\bm{\hat{G}}_{(i)}\big{)}^{\text{T}}$ and calculate the error $\epsilon_{(i)}=\big{\|}\big{[}\bm{\mathcal{Y}}\big{]}_{\left(3\right)}-\hat{\bm{Y}}_{(i)}\big{\|}_{\text{F}}^{2}$.

### v2

@@ IV Tensor signal modeling
In this section, we recast the signal model for the received pilots using a tensor decomposition approach. In correspondence with the background material presented in Section II, we start by formulating the tensor signal model for fully connected architectures. Then, we discuss the tensor signal formulation for the group-connected case, which is of practical interest due to its lower implementation complexity. The main expressions derived in this section will be exploited later to derive the proposed channel estimation methods.
IV-A Fully connected case
@@ IV-B Group-connected case
Let us now consider the group-connected architecture. Starting from the signal model in (14), and omitting the noise term, we have
@@
Group-connected architectures are commonly adopted due to implementation complexity. In this case, the BD-RIS training tensor is “sparse” due to its block-diagonal structure. Figure 2 illustrates the decomposition of received pilot signals in tensor form, corresponding to a Tucker-2 decomposition composed of $Q$ blocks. This decomposition can also be viewed as a special block-term decomposition (BTD), more specifically, also referred to as a “type-2 BTD”, which represents a tensor into a sum of rank-$(\bar{N},\bar{N},\cdot)$ tensor blocks [31, 36]. From this figure, we can see that the assumption of a group-connected architecture implies a BD-RIS training tensor having a “block-diagonal” structure, the $k$-th frontal slice of which corresponds to the BD-RIS training matrix associated with the $k$-th block. Note that a fully connected BD-RIS architecture is equivalent to having a fully dense BD-RIS training tensor in Figure 2, which does not have zero off-diagonal blocks. Although we focus on the group-connected case based on (28), any (sub-connected) architecture can be captured by the general tensor representation in (20), the difference being in the structure of the BD-RIS training tensor $\bm{\mathcal{S}}$.
Similarly to (24)-(26), we construct the unfoldings of the $q$-th BD-RIS scattering tensor as
@@
Hence, equivalent expressions can be obtained for the group-connected case for the 1-mode and 2-mode matrix unfoldings of the received pilot tensor, which can be obtained by rewriting (21) and (22) as a sum of $Q$ blocks
$\big{[}\bm{\mathcal{Y}}\big{]}_{\left(1\right)}=[\bm{G}^{(1)},\ldots,\bm{G}^{(Q)}]\cdot\left[\begin{array}[]{c}[\bm{\mathcal{S}}^{(1)}]_{(1)}(\bm{I}_{K}\otimes\bm{H}^{(1)\text{T}})\\ \vdots\\ [\bm{\mathcal{S}}^{(Q)}]_{(1)}(\bm{I}_{K}\otimes\bm{H}^{(Q)\text{T}})\end{array}\right]$ (35)
$\big{[}\bm{\mathcal{Y}}\big{]}_{\left(2\right)}=[\bm{H}^{(1)},\ldots,\bm{H}^{(Q)}]\cdot\left[\begin{array}[]{c}[\bm{\mathcal{S}}^{(1)}]_{(2)}(\bm{I}_{K}\otimes\bm{G}^{(1)\text{T}})\\ \vdots\\ [\bm{\mathcal{S}}^{(Q)}]_{(2)}(\bm{I}_{K}\otimes\bm{G}^{(Q)\text{T}})\end{array}\right]$ (39)
Likewise, the 3-mode matrix unfolding of the received pilot tensor can be obtained by rewriting (23) as a sum of $Q$ blocks
$\big{[}\bm{\mathcal{Y}}\big{]}_{\left(3\right)}$ $=$ $\Big{[}[\bm{\mathcal{S}}^{(1)}]_{(3)},\ldots,[\bm{\mathcal{S}}^{(Q)}]_{(3)}\Big{]}\left[\begin{array}[]{c}\bm{H}^{(1)\text{T}}\otimes\bm{G}^{(1)\text{T}}\\ \vdots\\ \bm{H}^{(Q)\text{T}}\otimes\bm{G}^{(Q)\text{T}}\end{array}\right]$ (43)
@@
Remark 1: One should note that the compact 3-mode unfolding expression in (45) is preferable to the one in (23) when considering a group-connected architecture. This is because the 3-mode unfolding of the BD-RIS tensor $\bm{\mathcal{S}}\in\mathbb{C}^{\bar{N}Q\times\bar{N}Q\times K}$ is sparse due to its block-diagonal structure (see Figure 2). On the other hand, the compact 3-mode unfolding $\bm{S}_{3}$ concatenates the 3-mode unfoldings of $Q$ tensor blocks. While $[\bm{\mathcal{S}}]_{(3)}$ is a sparse $K\times\bar{N}^{2}Q^{2}$ matrix with a $Q$ times factor increase on the number of columns, $\bm{S}_{3}$ is a dense $K\times\bar{N}^{2}Q$ matrix. Adopting the latter yields a more compact estimation of the composite channel in the group-connected case as the block Kronecker product of each group’s component. Hence, we adopt the compact 3-mode expression in (45) to formulate the block Tucker Kronecker factorization (BTKF) algorithm for channel estimation.
@@
which corresponds to a PARAFAC tensor model for the received pilots with single-connected RIS [23]. Table I summarizes signal models for different architectures: single-connected, group-connected, and fully connected, highlighting their differences, the resulting structure of the composite channel, and the related tensor model.
TABLE I: Comparison of signal models for different architectures: single-connected, group-connected, and fully connected.
Received signal (noise-free)
Composite channel
Connection degree ($N=\bar{N}Q$) Tensor model
$\bm{Y}_{k}=\bm{G}\text{diag}_{k}(\bm{S})\bm{H}^{\text{T}}$ $\bm{H}\diamond\bm{G}\in\mathbb{C}^{M_{R}M_{T}\times N}$ Single-connected $\bar{N}=1$, $Q=N$ PARAFAC [23]
$\bm{Y}_{k}=\sum\limits_{q=1}^{Q}\bm{G}^{(q)}\bm{S}_{k}^{(q)}\bm{H}^{(q)\text{T}}$ $\bm{H}\hskip 1.42271pt|\otimes|\hskip 1.42271pt\bm{G}\in\mathbb{C}^{M_{R}M_{T}\times\bar{N}^{2}Q}$ Group-connected, $\bar{N}\geq 2,Q\geq 2$ Block Tucker-$2$
$\bm{Y}_{k}=\bm{G}\bm{S}_{k}\bm{H}^{\text{T}}$ $\bm{H}\otimes\bm{G}\in\mathbb{C}^{M_{R}M_{T}\times N^{2}}$ Fully-connected $\bar{N}=N,Q=1$ Tucker-$2$
@@ V Tensor-Based channel estimation methods
In this section, we formulate the proposed channel estimation methods by capitalizing on the tensor signal structures discussed in the previous section. Two algorithms are proposed to solve the channel estimation problem via decoupling the estimates of the involved channel matrices $\bm{G}$ and $\bm{H}$. The first one, BTKF, is a closed-form solution that extracts the channel estimates via a block Kronecker factorization problem, which boils down to solving a set of rank-one matrix approximation problems. The second is BTALS, which directly estimates the channel matrices using an iterative procedure. We also discuss identifiability and uniqueness issues and their overall implications for system design. The design of the BD-RIS training tensor is a separate topic that will be addressed in the next section.
@@ V-A Block Tucker Kronecker factorization algorithm (BTKF) algorithm for decoupled channel estimation
where $\overline{\bm{Z}}^{(q)}\in\mathbb{C}^{M_{R}\bar{N}\times M_{T}\bar{N}}$ is a reshaped version of $\bm{Z}^{(q)}\in\mathbb{C}^{M_{R}M_{T}\times\bar{N}^{2}}$, $\bm{g}^{(q)}=\text{vec}\big{(}\bm{G}^{(q)}\big{)}\in\mathbb{C}^{M_{R}\bar{N}\times 1}$, and $\bm{h}^{(q)}=\text{vec}\big{(}\bm{H}^{(q)}\big{)}\in\mathbb{C}^{M_{T}\bar{N}\times 1}$. Please note that a rank-one matrix approximation is unique to a scaling factor affecting the estimated vectors. Thus, defining $\overline{\bm{Z}}^{(q)}=\bm{U}^{(q)}\bm{\Sigma}^{(q)}\bm{V}^{(q)\text{H}}$ the estimates of $\bm{g}^{(q)}$ and $\bm{h}^{(q)}$ correspond respectively to the dominant left and right singular vectors $\bm{u}^{(q)}_{1}=\bm{U}^{(q)}_{.1}\in\mathbb{C}^{M_{R}\bar{N}\times 1}$ and $\bm{v}^{*(q)}_{1}=\bm{V}^{*(q)}_{.1}\in\mathbb{C}^{M_{T}\bar{N}\times 1}$ associated with the first singular value $\sigma_{1}^{(q)}=\bm{\Sigma}^{(q)}_{1,1}$, i.e.,
$\hat{\bm{g}}^{(q)}$ $=\sqrt{\sigma_{1}^{(q)}}\bm{u}^{(q)}_{1},\quad\hat{\bm{G}}^{(q)}=\text{unvec}\big{(}\hat{\bm{g}}^{(q)}\big{)}_{M_{R}\times\bar{N}},$ (52)
$\hat{\bm{h}}^{(q)}$ $=\sqrt{\sigma_{1}^{(q)}}\bm{v}^{*(q)}_{1},\quad\hat{\bm{H}}^{(q)}=\text{unvec}\big{(}\hat{\bm{h}}^{(q)}\big{)}_{M_{T}\times\bar{N}}."$ (53)
@@ V-B Block Tucker alternating least squares (BTALS) algorithm for decoupled channel estimation
This problem is nonlinear in the unknowns since the solution involves products of the coefficients of the channel matrices $\bm{H}$ and $\bm{G}$. However, from tensor decomposition theory [38], this problem can be efficiently solved through an alternating least squares (ALS)estimation algorithm [38, 39], a popular iterative method for fitting a tensor model thanks to its simplicity and monotonic convergence property. In our case, the algorithm yields decoupled estimates of the channel matrices by converting the problem in (56) into two linear LS sub-problems that are solved iteratively.
The following developments consider the group-connected case due to its practical interest. [^3] The final expressions for the fully connected case can be easily deduced from the group-connected one by simply assuming $Q=1$ and $\bar{N}=N$. For later convenience, let us rewrite the expressions in (35)-(35) in compact form as
$\big{[}\bm{\mathcal{Y}}\big{]}_{\left(1\right)}=\bm{G}\bm{S}_{1}\bm{T}_{\bm{H}},\quad\big{[}\bm{\mathcal{Y}}\big{]}_{\left(2\right)}=\bm{H}\bm{S}_{2}\bm{T}_{\bm{G}},$ (57)
where $\bm{S}_{n}\doteq\text{blkdiag}\big{(}[\bm{\mathcal{S}}^{(1)}]_{(n)},\ldots,[\bm{\mathcal{S}}^{(Q)}]_{(n)}\big{)}\in\mathbb{C}^{\bar{NQ}\times\bar{N}KQ}$ collects the $n$-mode unfoldings of the BD-RIS tensor, $n=1,2$, while $\bm{T}_{\bm{H}}$ and $\bm{T}_{\bm{G}}$ are the Kronecker structured matrices
$\bm{T}_{\bm{H}}\doteq\big{[}\bm{I}_{K}\otimes\bm{H}^{(1)},\ldots,\bm{I}_{K}\otimes\bm{H}^{(Q)}\big{]}^{\text{T}}\in\mathbb{C}^{\bar{N}KQ\times M_{T}K},$ (58)
$\bm{T}_{\bm{G}}\doteq\big{[}\bm{I}_{K}\otimes\bm{G}^{(1)},\ldots,\bm{I}_{K}\otimes\bm{G}^{(Q)}\big{]}^{\text{T}}\in\mathbb{C}^{\bar{N}KQ\times M_{R}K}.$ (59)
These expressions are exploited to directly estimate $\bm{G}$ and $\bm{H}$ by iteratively solving the following two LS problems
$\hat{\bm{G}}$ $=\underset{\bm{G}}{\mathop{\rm argmin}}\left\|\big{[}\bm{\mathcal{Y}}\big{]}_{\left(1\right)}-\bm{G}\bm{S}_{1}\bm{T}_{\bm{H}}\right\|_{\rm F}^{2},$ (60)
$\hat{\bm{H}}$ $=\underset{\bm{H}}{\mathop{\rm argmin}}\left\|\big{[}\bm{\mathcal{Y}}\big{]}_{\left(2\right)}-\bm{H}\bm{S}_{2}\bm{T}_{\bm{G}}\right\|_{\rm F}^{2},$ (61)
@@
$\hat{\bm{G}}=\big{[}\bm{\mathcal{Y}}\big{]}_{\left(1\right)}\left[\bm{S}_{1}\bm{T}_{\bm{H}}\right]^{+},\quad\hat{\bm{H}}=\big{[}\bm{\mathcal{Y}}\big{]}_{\left(2\right)}\left[\bm{S}_{2}\bm{T}_{\bm{G}}\right]^{+}.$ (62)
The algorithm intertwines the LS estimates of $\bm{G}$ and $\bm{H}$. At each iteration, one matrix is updated based on a previously obtained estimate of the other matrix. This procedure is repeated until convergence. An error measure computed at the end of the $i$-th iteration is given by
$\epsilon_{(i)}\doteq\frac{\left\|\big{[}\bm{\mathcal{Y}}\big{]}_{\left(3\right)}-\hat{\bm{Y}}_{(i)}\right\|_{\rm F}^{2}}{\left\|\big{[}\bm{\mathcal{Y}}\big{]}_{\left(3\right)}\right\|_{\rm F}^{2}},$ (63)
where $\hat{\bm{Y}}_{(i)}\doteq[\bm{\mathcal{S}}]_{(3)}\big{(}\bm{\hat{H}}_{(i)}\otimes\bm{\hat{G}}_{(i)}\big{)}^{\text{T}}$ is the reconstructed signal model at the $i$-th iteration. The convergence is declared when the error does not significantly change between two successive iterations, which implies that $|\epsilon_{(i)}-\epsilon_{(i-1)}|\leq\eta$, where $\eta=10^{-6}$ is a convergence threshold. A summary of the steps of the proposed BALS algorithm is provided in Algorithm 2.
Such an alternating estimation procedure may improve or maintain but cannot worsen the current fit, usually leading to a global minimum solution [14]. For simplicity, random initialization can be adopted. However, one may resort to enhanced initialization and acceleration procedures [14, 38]. Other initialization strategies that could be used to exploit the structure of the channel matrices are possible, although not advocated here. In our case, however, where the core tensor (represented by the BD-RIS structure) is known at the receiver, such enhancements are not required, and random initialization is enough to obtain satisfactory results.
@@
6:     Define the singular value decomposition (SVD) of $\overline{\bm{Z}}^{(q)}=\bm{U}^{(q)}\bm{\Sigma}^{(q)}\bm{V}^{(q)\text{H}}$.
@@
$\hat{\bm{G}}^{(q)}$ $=\text{unvec}\big{(}\sqrt{\sigma_{1}^{(q)}}\bm{u}^{(q)}_{1}\big{)}_{M_{R}\times\bar{N}},$
$\hat{\bm{H}}^{(q)}$ $=\text{unvec}\big{(}\sqrt{\sigma_{1}^{(q)}}\bm{v}^{*(q)}_{1}\big{)}_{M_{T}\times\bar{N}}.$
@@
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bm{T}_{\hat{\bm{H}}(i-1)}\doteq\Big{[}\bm{I}_{K}\otimes\bm{H}^{(1)}_{(i-1)},\ldots,\bm{I}_{K}\otimes\bm{H}^{(Q)}_{(i-1)}\Big{]}^{\text{T}}}$
$\bm{\hat{G}}_{(i)}=\big{[}\bm{\mathcal{Y}}\big{]}_{\left(1\right)}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\left[\bm{S}_{1}\bm{T}_{\hat{\bm{H}}(i-1)}\right]^{+}}.$
@@
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bm{T}_{\hat{\bm{G}}(i)}\doteq\Big{[}\bm{I}_{K}\otimes\bm{G}^{(1)}_{(i)},\ldots,\bm{I}_{K}\otimes\bm{G}^{(Q)}_{(i)}\Big{]}^{\text{T}}}$
$\bm{\hat{H}}_{(i)}=\big{[}\bm{\mathcal{Y}}\big{]}_{\left(2\right)}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\left[\bm{S}_{2}\bm{T}_{\hat{\bm{G}}_{(i)}}\right]^{+}}.$
6:     Compute $\hat{\bm{Y}}{(i)}=[\bm{\mathcal{S}}]_{(3)}\big{(}\bm{\hat{H}}_{(i)}\otimes\bm{\hat{G}}_{(i)}\big{)}^{\text{T}}$ and calculate the error according to (63).